OpenStudy (anonymous):

Use green's theorem to evaluate the line integral along the given positive orientated curve. \[\int_c xe^{-2x}dx+(x^4+2x^2y^2)dy\] C is the boundary of the region between the circles \[x^2+y^2=1\] and \[x^2+y^2=4\] \[\int\int_D\left[\frac{\partial}{\partial x}(x^4+2x^2y^2)-\frac{\partial}{\partial y}(xe^{-2x})\right]dA \] \[\int\int_D(4x^3+4xy^2)dA\] would I do greens theorem again? \[\int\int_D\left[\frac{\partial}{\partial x}(4xy^2)-\frac{\partial}{\partial y}(4x^3)\right]dA \] \[\int\int_D4y^2dA\]

OpenStudy (anonymous):

to my understanding for last stage u just need to compute double integral

OpenStudy (turingtest):

yeah, and I would do it in polar coordinates

OpenStudy (turingtest):

\[\iint\limits_D(4x^3+4xy^2)dA\]\[D:\{(r,\theta)|1\le r\le2,0\le\theta\le2\pi\}\]\[x=r\cos\theta~~~~~~~y=r\sin\theta~~~~~~~~dA=rdrd\theta\]

OpenStudy (anonymous):

I understand everything you wrote except for \[dA=rdrd\theta\]

OpenStudy (turingtest):

that is dA in polar coordinates... if you want to prove that you need to use the jacobian (not sure if you're doing that yet)

OpenStudy (anonymous):

not yet

OpenStudy (turingtest):

so then you have to take my word for it that you need to add another r in the dA for polar coordinates or http://en.wikipedia.org/wiki/Polar_coordinate_system

OpenStudy (anonymous):

fair enough. So now we have: \[\int_1^2\int_0^{2\pi}(4x^3+4xy^2)rdrd\theta\]

OpenStudy (anonymous):

Nope that's wrong, let me try it again

OpenStudy (turingtest):

gotta sub in for x and y

OpenStudy (anonymous):

\[\int_1^2\int_0^{2\pi}(4(rcos\theta)^3+4(rcos\theta)(rsin\theta)^2)rdrd\theta\]

OpenStudy (anonymous):

\[\int_1^2\int_0^{2\pi}(4r^4cos^3+4r^4cos\theta sin^2\theta)drd\theta\]

OpenStudy (turingtest):

yep now use some identities

OpenStudy (turingtest):

well, one identity just factor it right...

OpenStudy (anonymous):

would I take the 4r^4 out of the first integral since we're doing the integral with respect to theta first?

OpenStudy (turingtest):

try taking out 4r^4cos(theta)

OpenStudy (turingtest):

I mean factoring it that way rather

OpenStudy (anonymous):

that would give me \[cos^2\theta+sin^2\theta=1\] and now i would take the integral of \[\int_1^2\int_0^{2\pi}4r^4cos\theta drd\theta\]

OpenStudy (turingtest):

yes, and you will see quickly what it becomes

OpenStudy (anonymous):

we would be left with \[\int_1^24r^4dr\] because the integral of cosine is sine and then (0-0) \[4\left[\frac{r^5}{5}\right]_1^2=4(\frac{32}{5}-\frac15)\]

OpenStudy (turingtest):

what is\[\int_0^{2\pi}\cos\theta d\theta\]?

OpenStudy (anonymous):

is sine

OpenStudy (turingtest):

and after evaluating...?

OpenStudy (anonymous):

sine of 2pi is zero and so is sine of zero

OpenStudy (turingtest):

and so that's where we end, because the integral of zero is zero so the next one goes away....

OpenStudy (anonymous):

oh that would make the double integral zero?

OpenStudy (turingtest):

yep

OpenStudy (turingtest):

I hate when that happens; always makes me think I messed up but here I don't think so

OpenStudy (anonymous):

oh ok. Thanks! Does that mean there is no area between the region bounded by those two circles?

OpenStudy (turingtest):

no, it means that when you integrate that function over the area the answer is zero. obviously the area between a circle radius 1 and radius 2 itself is not zero; it can easily be calculated with old-fashioned geometry

OpenStudy (turingtest):

which, according to green's theorem, means that the line integral is zero

OpenStudy (anonymous):

I thought we were integrating to find the area?

OpenStudy (anonymous):

oh sorry we're evaluating the curve

OpenStudy (turingtest):

no, and area integral would be\[\iint\limits_D dA\]and nothing more that is different than\[\iint\limits_Df(x)dA\]in which case you are integrating a \(function\) over an area

OpenStudy (turingtest):

greens theorem relates a line integral over a closed curve to a surface integral of a related function

OpenStudy (turingtest):

\[\oint\limits_C Pdx+Qdy=\iint\limits_D\left(\frac{\partial Q}{\partial x}-\frac{\partial P}{\partial y}\right)dA\]

OpenStudy (turingtest):

it's a relation between the two ideas; a closed line integral and the surface integral on the right.

OpenStudy (anonymous):

I think I'm starting to get the idea now. I'll do more problems to get a better grip on greens theorem.

OpenStudy (turingtest):

here's a good reference: http://tutorial.math.lamar.edu/Classes/CalcIII/GreensTheorem.aspx

OpenStudy (anonymous):

I can't believe I forgot to reference Paul again!

OpenStudy (turingtest):

yeah, apul has a similar problem in example 3 a note about LaTeX: to make the double integral look better, and to get the D underneath the signs, try \iint\limits_D\[\iint\limits_D\] likewise \iiint\[\iiint\] is a triple integral and \oint\[\oint\]is a line integral

OpenStudy (turingtest):

paul*

OpenStudy (anonymous):

Yeah i thought my integrals were quite far apart. I'll do that for this next problem i'm working on. Thanks buddy!

OpenStudy (turingtest):

\oint is a *closed* line integral I mean..

OpenStudy (turingtest):

anytime!

OpenStudy (anonymous):

dude I forgot to give you a medal!